OpenStudy (montsechamorro):

1) What is the equation of a circle that passes through (2, 0) and has its center at M(4, 0)? 2)What is the equation of a circle that passes through point J(6, 4) and has center K (1, -8).

OpenStudy (montsechamorro):

help<3

undeadknight26 (undeadknight26):

IM trying Im trying.

OpenStudy (montsechamorro):

oh thank you so much!

TheSmartOne (thesmartone):

use the distance formula to calculate the distance between the two points the distance is the radius do you know the distance formula?

OpenStudy (montsechamorro):

no i don't ;'(

TheSmartOne (thesmartone):

http://prntscr.com/bzcytm

OpenStudy (montsechamorro):

so the distance is 2?

OpenStudy (montsechamorro):

for #1?

TheSmartOne (thesmartone):

correct, that means the radius is 2

OpenStudy (montsechamorro):

okay so then what do i do

TheSmartOne (thesmartone):

now the equation of a circle is (x - h)^2 + (y- k)^2 = r^2 where (h, k) is the center and r is the radius

OpenStudy (montsechamorro):

so h is 4? and k is 0?

TheSmartOne (thesmartone):

correct :)

OpenStudy (montsechamorro):

so my formula is (2-4)^2+(0-0)^2=4

OpenStudy (montsechamorro):

oh wait!

TheSmartOne (thesmartone):

you don't replace x with 2 and y with 0 :P

OpenStudy (montsechamorro):

so i just fill in k and h

TheSmartOne (thesmartone):

and r

OpenStudy (montsechamorro):

gotcha so my answer would be, (x-4)^2+y^2=4

TheSmartOne (thesmartone):

you're using the x-value of one point and the y-value of another point you have to be consistent, and you need to use the center's point

OpenStudy (montsechamorro):

Im confused

OpenStudy (montsechamorro):

wouldn't that just be the equation for #1?

TheSmartOne (thesmartone):

sorry, it's really late here and I don't know what I read when you posted that earlier your answer is correct but I think you should leave it as (y - 0)^2

OpenStudy (montsechamorro):

alrighty thank you, I did #2 but i just wanna make sure its correct, i got (x-1)^2+(y+8)^2=169

TheSmartOne (thesmartone):

Correct! Great work! :D

OpenStudy (montsechamorro):

I have one more problem i just don't know how to start it!

TheSmartOne (thesmartone):

Sure, post ahead and I'll see if I can help :)

OpenStudy (montsechamorro):

What is the equation of this graph? (

TheSmartOne (thesmartone):

well, it's increasing by 2 on the x-axis and y-axis the center is at (2, -1) and a point that the circle passes through is (8, 0)

OpenStudy (montsechamorro):

hmm i got my distance as 6.08

OpenStudy (montsechamorro):

should i just leave it at 6?

OpenStudy (montsechamorro):

i did it, and i got (x-2)^2+(y+1)^2=36

TheSmartOne (thesmartone):

keep it as sqrt(37) because that is the radius, and the equation wants radius^2 which would make it (sqrt(37))^2 = 37

OpenStudy (montsechamorro):

Okay thank you so much! I got a 100% thanks to your help:)

TheSmartOne (thesmartone):

Anytime! :D

TheSmartOne (thesmartone):

Thank you for cooperating and learning! ^-^

OpenStudy (montsechamorro):

have a good night!

TheSmartOne (thesmartone):

You too! :D